OpenStudy (anonymous):

find the angle between the following vectors: A=12i-15j+25k , b=-12i+20j+30k please show me steps and explanation :)

OpenStudy (amistre64):

if you know the magnitude of each angle, then its just an application of the law of cosines

OpenStudy (amistre64):

... for the law of cosines in this one, you would also need to know the length (magnitude) of the resultant vector from the sum of the other 2

OpenStudy (anonymous):

snx :)

OpenStudy (amistre64):

a= 12i-15j+25k + b=-12i+20j+30k ------------------ c = 0i + 5j +55k

OpenStudy (amistre64):

once the lengths of each vector is known; the law of cosines can be constructed as:\[c^2=a^2+b^2-2ab~cos(\theta)\] \[c^2-a^2-b^2=-2ab~cos(\theta)\] \[\frac{c^2-a^2-b^2}{-2ab}=cos(\theta)\] \[cos^{-1}\left(\frac{c^2-a^2-b^2}{-2ab}\right)=\theta\]

OpenStudy (amistre64):

theres also a dot product version of this :)

OpenStudy (amistre64):

\[cos~\theta=\frac{a.b}{|a|~|b|}\] \[\theta=cos^{-1}\left(\frac{a.b}{|a|~|b|}\right)\] and that is proved by the law of cosines

OpenStudy (anonymous):

theta=1.3125